# EEdit [Uncaptioned image] : Rethinking the Spatial and Temporal Redundancy for Efficient Image Editing

### v2

@@ Abstract
Inversion-based image editing is rapidly gaining momentum while suffering from significant computation overhead, hindering its application in real-time interactive scenarios. In this paper, we rethink that the redundancy in inversion-based image editing exists in both the spatial and temporal dimensions, such as the unnecessary computation in unedited regions and the redundancy in the inversion progress. To tackle these challenges, we propose a practical framework, named EEdit, to achieve efficient image editing. Specifically, we introduce three techniques to solve them one by one. For spatial redundancy, spatial locality caching is introduced to compute the edited region and its neighboring regions while skipping the unedited regions, and token indexing preprocessing is designed to further accelerate the caching. For temporal redundancy, inversion step skipping is proposed to reuse the latent for efficient editing. Our experiments demonstrate an average of 2.46$\times$ acceleration without performance drop in a wide range of editing tasks including prompt-guided image editing, dragging and image composition. Our codes are available at https://github.com/yuriYanZeXuan/EEdit.
@@ 1 Introduction
We validated the effectiveness of proposed modules in extensive experiments on various editing tasks, including prompt-guided [49, 7, 15], reference-guided [26, 53, 50], and drag-guided editing [36, 24, 55, 42]. While achieving state-of-the-art performance in background consistency, EEdit also delivers up to 10.96$\times$ latency acceleration compared to other editing methods. Even when compared to existing cache-based acceleration techniques, it demonstrates a superior acceleration ratio. Furthermore, EEdit achieves a 2.46$\times$ speedup with minimal performance differences, maintaining near-lossless quality across various evaluation metrics when compared to the original editing pipeline without cache acceleration.
@@
  • Extensive Adaption and Experiments: We evaluate our methods across various editing tasks, including prompt-guided editing, drag-guided editing, and reference-guided editing, which demonstrates 10.96$\times$ acceleration than the current state-of-the-art editing approach.
Figure 3: The overview of our approach. The proposed framework for image editing based on MM-DiT diffusion models employs an efficient denoising and training-free approach. The pipeline takes the original image and an editing prompt as inputs. Specifically, the cache is refreshed entirely in fixed time-step interval, while partial computation for updating cache is maintained for the intermediate timesteps.
@@ 2.1 Acceleration of Diffusion Models
First, computing intrinsic feature correlations between tokens introduces additional internal overhead in caching operations. Second, some caching strategies and editing methods require accessing attention maps [59] or storing KV matrices [26, 58], leading to incompatibility with existing Transformer acceleration techniques, such as FlashAttention [8], thereby increasing latency. Moreover, editing tasks inherently exhibit prior information regarding the spatial distribution of regions requiring greater attention, yet current caching techniques fail to leverage this information.
@@ 3.1 Rectified Flow
[…]
@@ 4 Approaches
4:$\mathcal{M}_{freq}\leftarrow\boldsymbol{zero}\mathbf{[\mathcal{F}_{i},l]}$
@@
12:$\mathbf{Z}_{0}\leftarrow\mathbf{cat}(\text{VQ-Encoder}(\mathbf{I_{s}}),\text{Txt-Encoder}(\mathbf{P_{m}}))$
@@ 4.1 Spatial Locality Caching
Here $\mathbf{x}$ represents an arbitrary token in DiT. $M_{s}$ is the latent mask for point set of edit region. Hyperparameters such as $b>1$ represent the bonus factor, while $10<r<$ 1 denotes the decay ratio. $K$ is the maximum number of layers for the neighborhood. The neighborhood of $M_{s}$ with an L1 norm equal to $k$ is denoted as $\mathcal{N}_{k}(M_{s})$ and is defined as:
@@ 4.2 Token Index Preprocessing
[…]
@@ 5.2 Compare with Baseline
Qualitative Comparison. We perform extensive qualitative comparisons between our method and various editing approaches. Specifically, we evaluate three kinds of popular editing tasks, including prompt-guided, drag-guided, and reference-guided image editing. (1) Prompt-guided image editing: we compare with Direct Inversion  [15], InfEdit [52], and P2P [13]. (2) Drag-guided image editing, we select three methods to do comparisons, including DragDiffusion [42], Region Drag[24], and FastDrag [55]. Pipeline are implemented from public codebase from github. (3) Reference-guided editing, we compare our approach with TF-ICON [26], PrimeComposer [50], and DCCF [53]. SD-v2-1 checkpoint is used for Image Composition. As shown in Figure. 5, for prompt-guided task, Direct inversion and InfEdit fail to edit the image successfluly. DDIM$+$P2P has a challenge to maintain the background consistency. By contract, our approach exhibits superior consistency, enhanced detail preservation, and improved aesthetic quality in terms of style.
Table 1: Comparison of quality across different editing methods.
Editing methods Inversion type PSNR $\uparrow$ LPIPS ×${}^{\downarrow}_{\times 10^{-2}}$ SSIM $\uparrow$ CLIP-T $\uparrow$
P2P [13] DDIM 17.87 20.88 0.72 25.13
MasaCtrl [4] DDIM 22.19 10.54 0.80 24.02
DI 22.69 8.73 0.82 24.39
PnP [15] DDIM 25.23 11.27 0.80 25.42
DI 22.46 10.55 0.80 25.48
InfEdit [52] Inversion-free 28.11 5.61 0.85 25.86
DiT4Edit  [10] DPM-Solver++ 22.85 - - 25.39
SLoC RF-inversion 31.97 1.96 0.94 25.37
SLoC ISS 31.97 1.95 0.94 25.38
@@ 5.3 Ablation Study
Ablation study of TIP over ISS and various tasks. As shown in Table 3 and Figure 8, TIP remains compatible with ISS while maintaining comparable FG fidelity and CLIP scores across various editing tasks. Furthermore, it achieves an additional reduction in inference latency by an average of over 20% compared to the SLoC baseline. Under the same configuration, the impact of incorporating TIP on editing quality remains within the range of random fluctuations. Furthermore, our quantitative experiments validate that this approach achieves lossless acceleration while maintaining the same generation quality which can be seen in supplementary material.
@@ 8.2 Vanilla SLoC w/o ISS and TIP
3:$\mathcal{M}_{freq}\leftarrow\boldsymbol{zero}\mathbf{[\mathcal{F}_{i},t,l]}$
4:$\mathbf{Z}_{0}\leftarrow\mathbf{cat}(\text{VQ-Encoder}(\mathbf{I_{s}}),\text{Txt-Encoder}(\mathbf{P_{m}}))$
@@
[…]
@@ 9.2 Notation and Problem Setup
  • $N$: Total number of tokens, indexed as $\{1,2,\dots,N\}$.
@@ 9.5 Proof of Equivalence
10 Implementation Details
The experiments were conducted on a machine with the following hardware and software specifications:
10.1 Hardware Specifications
  • Architecture: x86_64
  • CPU Op-Modes: 32-bit, 64-bit
  • Address Sizes: 52 bits physical, 48 bits virtual
  • Byte Order: Little Endian
  • Total CPU(s): 128
  • On-line CPU(s) List: 0-127
  • Vendor ID: AuthenticAMD
  • Model Name: AMD EPYC 9K84 96-Core Processor
  • CPU Family: 25
10.2 Software Specifications
  • Operating System: Ubuntu 22.04.3 LTS
  • Python 3.12.3
  • huggingface-hub: 0.26.2
  • numpy: 1.26.4
  • torch: 2.5.1
  • torchmetrics: 1.6.1
  • transformers: 4.46.1
@@ 12 More Experiments
We compared the edited results in terms of image similarity and efficiency. The Table. 4 demonstrate that our approach, when incorporating all optimizations (TIP + ISS), achieves the best performance in both fidelity and efficiency. SLoC achieves a 2.46$\times$ significant improvement in inference latency and a 3.53$\times$ acceleration in computational efficiency compared to the original unaccelerated version. Additionally, our method demonstrates either improved fidelity or remains at a state-of-the-art level across various editing tasks.

### v3

@@ Abstract
[…]
@@ 1 Introduction
[…]
@@
[…]
Figure 3: The overview of our approach. The proposed framework for image editing based on $M$-DiT diffusion models employs an efficient denoising and training-free approach. The pipeline takes the original image and an editing prompt as inputs. Specifically, the cache is refreshed entirely in fixed time-step interval, while partial computation for updating cache is maintained for the intermediate timesteps.
@@ 2.1 Acceleration of Diffusion Models
First, computing intrinsic feature correlations between tokens introduces additional internal overhead in caching operations. Second, some caching strategies and editing methods require accessing attention maps [105] or storing KV matrices [39, 104], leading to incompatibility with existing Transformer acceleration techniques, such as FlashAttention [11, 92], thereby increasing latency. Moreover, editing tasks inherently exhibit prior information regarding the spatial distribution of regions requiring greater attention, yet current caching techniques fail to leverage it.
@@ 3.1 Rectified Flow
$\min_{\theta}\int_{0}^{1}\mathbb{E}\left[\left\|\mathbf{X}_{1}-\mathbf{X}_{0}-v_{\theta}(\mathbf{X}_{t},t)\right\|^{2}\right]dt$
@@ 4 Approaches
1:Input image $\mathbf{I_{s}}$, Mask for editing region $\mathbf{M_{s}}$, Prompt for editing $\mathbf{P_{m}}$, Randomly initialized map $\mathcal{R}$, Bonus for edited region $\mathbf{S_{E}}$, Cache dict $\mathcal{C}_{t}[l,f]$ and NN layer $\mathcal{F}_{i}$. Token-wise multiplication $\odot$ and addition $\oplus$.
@@
[…]
@@
[…]
@@ 4.1 Spatial Locality Caching
Here $\mathbf{x}$ represents an arbitrary token in DiT. $M_{s}$ is the latent mask for point set of edit region. Hyperparameters such as $b>1$ represent the bonus factor, while $0<r<1$ denotes the decay ratio. $K$ is the border‐score smoothing hyperparameter. The neighborhood of $M_{s}$ with an L1 norm equal to $k$ is denoted as $\mathcal{N}_{k}(M_{s})$ and is defined as:
@@ 4.2 Token Index Preprocessing
we can prove the top-$R\%$ selected indices $\mathcal{I}^{(t)}_{\text{topR\%}}$ in the same time step in the offline and online process remain equivalent:
$\mathcal{I}^{(t)}_{\text{topR\%}}(\text{offline})=\mathcal{I}^{(t)}_{\text{topR\%}}(\text{online})\quad\forall t\in[1\dots T].$
@@ 5.2 Compare with Baseline
Qualitative Comparison. We perform extensive qualitative comparisons between our method and various editing approaches. Specifically, we evaluate three kinds of popular editing tasks, including prompt-guided, drag-guided, and reference-guided image editing. (1) Prompt-guided image editing: we compare with Direct Inversion  [26], InfEdit [87], and P2P [21]. (2) Drag-guided image editing, we select three methods to do comparisons, including DragDiffusion [63], Region Drag[36], and FastDrag [99]. Pipeline are implemented from public codebase from github. (3) Reference-guided editing, we compare our approach with TF-ICON [39], PrimeComposer [82], and DCCF [88]. SD-v2-1 checkpoint is used for Image Composition. As shown in Figure. 5, for prompt-guided task, Direct inversion and InfEdit fail to edit the image successfluly. DDIM++P2P has a challenge to maintain the background consistency. By contrast, our approach exhibits superior consistency, enhanced detail preservation, and improved aesthetic quality in terms of style.
Table 1: Comparison of quality and efficiency
Method Inversion Quality metrics Efficiency metrics
[…]
P2P DDIM 17.87 20.88 0.72 25.13 334.4 18.75 SD1-4 860M
MasaCtrl DDIM 22.19 10.54 0.80 24.02 334.4 17.65 SD1-4 860M
MasaCtrl DI 22.69 8.73 0.82 24.39 408.3 23.60 SD1-4 860M
PnP DDIM 25.23 11.27 0.80 25.42 180.9 13.81 SD1-5 860M
PnP DI 22.46 10.55 0.80 25.48 180.9 13.77 SD1-5 860M
InfEdit Inversion-free 28.11 5.61 0.85 25.86 124.6 2.90 SD1-5, LCM 860M
DiT4Edit DPM-Solver++ 22.85 25.39 260.4 5.15 PixArt XL 610M
RF-inv RF-inversion 17.74 24.40 0.66 26.31 1111.6 13.56 FLUX 12B
RF-Edit RF-Solver 20.17 18.50 0.77 26.64 2223.2 26.23 FLUX 12B
Flow-Edit 22.20 10.49 0.85 25.80 952.8 11.84 FLUX 12B
SLoC RF-inversion 31.97 1.96 0.94 25.37 384.0 5.96 FLUX 12B
SLoC ISS 31.97 1.95 0.94 25.38 264.5 4.60 FLUX 12B
@@ 5.3 Ablation Study
Ablation study of TIP over ISS and various tasks. As shown in Table 3 and Figure 8, TIP remains compatible with ISS while maintaining comparable FG fidelity and CLIP scores across various editing tasks. Furthermore, it achieves an additional reduction in inference latency by an average of over 20% compared to the SLoC baseline.
@@ 6 Conclusions
Acknowledgment This work was partially supported by Dream Set Off - Kunpeng&Ascend Seed Program and The Research Grant Council of the Hong Kong Special Administrative Region under grant number 16203122.
@@ 8.2 Vanilla SLoC w/o ISS and TIP
[…]
@@
SLoC+TIP+ISS 39.21 31.75 1.71 9.48 28 1.92 2.49 4.60 264.50 $\mathbf{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\textbf{2.46}\times}}$ $\mathbf{{\color[rgb]{0.1328125,0.546875,0.1328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1328125,0.546875,0.1328125}\textbf{3.53}\times}}$
@@ 9.2 Notation and Problem Setup
  • NN: Total number of tokens, indexed as $\{1,2,\dots,N\}$.
@@ 11 More Experiments
[…]
12 Related work
We thanks the related work about editing both image and video, such as VideoGrain [92], FastVAR [19], IntLora [18], FollowFamilty [47, 50, 42, 44, 49, 52, 90, 96, 103, 79, 14, 5], Tpsence [101], Pointnorm [100], MotionDiff [51], FastScene [48], DreamRelation [85], Dreamvideo [84], and Lazymar [89] and so on  [6, 66, 68, 67].
@@ 12 Related work
12.1 AI safety discussion
As discussed in previous works such as watermark techniques [75, 71, 24, 7, 8, 33, 91], AI-generated content should be more responsible and safe by adding watermark sign. We will release relative code to make sure our work be safe and not serve for harmful circumstance.
12.2 More relative controllable methods
In more recent works, controllable generation of diffusion models remains a popular topic, such as  [97, 94, 72, 73, 23, 74, 20, 70, 69, 98, 95, 78, 83, 16, 25, 37, 61, 62] These works inspire us to apply locally sensitive caching in a wider range of controllable generation and editing tasks to achieve faster generation acceleration.